OpenStudy (anonymous):

Solve 2x^2 - 5x + 1 = 3 for x.

OpenStudy (anonymous):

I think you will need to use quadratic equation or completing square. Below is the quadratic equation: \[x = \frac{-b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] for your question 2x^2 - 5x + 1 = 3 2x^2 - 5x - 2 = 0 So a = 2, b = -5, c = -2 Therefore, you just substitute the a, b and c into the equation: \[x = \frac{-(-5) \pm \sqrt{(-5) ^{2} - 4(2)(-2)} }{ 2(2)}\] \[x = \frac{5 \pm \sqrt{25+16} }{ 4}\] \[x = \frac{5 \pm \sqrt{41} }{ 4}\]

jhonyy9 (jhonyy9):

nice @Orca !!!

OpenStudy (anonymous):

Thanks! :) @jhonyy9